OpenStudy (anonymous):

which property justifies th statement 2 + (8 + 9) = (2 + 8) + 9 transitive commutative distributive associative

OpenStudy (anonymous):

@DROPD33DM9

OpenStudy (anonymous):

@brandonp7

OpenStudy (anonymous):

@razor99

OpenStudy (anonymous):

@jacobrocket173

OpenStudy (anonymous):

@paulalovee

OpenStudy (anonymous):

@kohai

OpenStudy (anonymous):

if im correct distributive: a(b+c)=ab+ac associative: a+b=b+a cumulative: a+(b+c)=(a+b)+c

OpenStudy (anonymous):

commutative*

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

and transitive is if a=b and b=c the a=c